OpenStudy (anonymous):

If n is even, which of the following cannot be odd? I n + 3 II 3n III n² - 1

OpenStudy (zzr0ck3r):

let n = 2 2+3 = 5 3*2 = 6 2^2-1 = 3

OpenStudy (zzr0ck3r):

can you answer now?

OpenStudy (zzr0ck3r):

if n is even 3 is always odd

OpenStudy (zzr0ck3r):

you can do this in a more general way if n is even the n = 2k for some integer k n+3 = 2k+3 is always odd 3n = 3*2k = 6k is always even n^2-1=(2n)^2-1 = 4n^2-1 is always odd